OpenStudy (unklerhaukus):

\[\int\frac{\text d v}{\sqrt{1+v^2}}\]

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

integrate

OpenStudy (anonymous):

look in the back of the text. your integrand is the derivative of a famous function

OpenStudy (anonymous):

in other words there are no steps here, no substitutions, no parts, no nothing

OpenStudy (unklerhaukus):

i dont want \[=\sinh^{-1}(x)+c\]

OpenStudy (anonymous):

what would you prefer?

OpenStudy (anonymous):

i mean it is going to be that in any case

OpenStudy (anonymous):

you could make \(v=\tan(x)\) to get rid of the radical

OpenStudy (anonymous):

get \[\int \frac{sec^2(x)dx}{\sqrt{\tan^2(x)+1}}\] \[\int \sec(x)dx\] etc

OpenStudy (unklerhaukus):

i want to show it \[=\ln \left(v+\sqrt{1+v^2}\right)\]

OpenStudy (anonymous):

not sure why though ok then this should work

OpenStudy (anonymous):

anti derivative of secant is \(\ln(\sec(x)+\tan(x)\) substitute back and you have your answer

OpenStudy (anonymous):

will that work out ok?

OpenStudy (mimi_x3):

\[\int\limits\frac{1}{\sqrt{1+v^{2}}} dx\] have you tried \(1+v^2=y^2\) => \(2xdx=2ydy\)

OpenStudy (anonymous):

http://math2.org/math/integrals/more/sec.htm i assume you 've probably forgotten the integration of secx

OpenStudy (unklerhaukus):

i dont remember ever doing that @rs32623

OpenStudy (unklerhaukus):

I couldn't get either of those methods to the end fortunately integral 25 in the back of my book has \[\int\frac{1}{\sqrt{x^2\pm a^2}}=\ln|x+\sqrt{x^2\pm a^2}| +c\]

OpenStudy (anonymous):

90% of what you learn in calc 2 is in the back two pages of the book, which is why that is my least favorite subject in all of math

OpenStudy (anonymous):

i don't see a point why can't we reach the result by putting the value of secx & tanx in the method described by satelite73 respectively (1+v)^2 & v..

OpenStudy (unklerhaukus):

i tried that

OpenStudy (anonymous):

we can, but it hinges on knowing that \(\int \sec(x)dx=\ln(\sec(x)+\tan(x))\)

OpenStudy (anonymous):

oh we can do that easily

OpenStudy (unklerhaukus):

where does that come from

OpenStudy (anonymous):

@UnkleRhaukus i've already provided the link for that in my first reply..

OpenStudy (unklerhaukus):

sorry @rs32623 i read that page i just didn't understand it very well

OpenStudy (anonymous):

if you grant that \[\int\sec(x)dx=\ln(\sec(x)+\tan(x))\] substitute back. since \(v=\sec(x)\iff x=\sec^{-1}(x)\) we know \[tan(x)=\tan(\sec^{-1}(x))=\sqrt{1+v^2}\] substitute back and get \[\ln(v+\sqrt{1+v^2})\] as required

OpenStudy (anonymous):

the tricky part is finding \(\int \sec(x)dx\)

OpenStudy (anonymous):

@UnkleRhaukus is the substituting back part ok? that is a set required in all trig substitutions

OpenStudy (anonymous):

*step

OpenStudy (anonymous):

where did you miss? it was pretty simple..starting from multiplying & dividing by secx+tanx & assuming secx+tanx=u & differentiating both side wrt x & u respectively we have secx(secx+tanx)dx=du so that the last terms is du/u..

OpenStudy (anonymous):

as for finding the anti derivative of secant, there is a rather stupid gimmick you can use, namely multiplying top and bottom by \(\sec(x)+tan(x)\) but if you haven't seen it would not know it. there is a more (to my mind) satisfactory method, which is to multiply top and bottom by \(\cos(x)\) replace \(cos^2(x)\) in the denominator by \(1-\sin^2(x)\) and then use partial fractions. both are rather gimmicky, but that is the nature of finding anti derivative s

OpenStudy (anonymous):

that is why, as i said. 90% of what you need in calc 2 is in the back two pages of the book, and frankly now that we have computer algebra systems like maple and mathematica (wolfram) we don't even need those two pages. a collection of show offy stupid gimmicks, so professors can say look i can integrate this look i can integrate that when the truth of the matter is that if you write down a function the probability it has a nice closed form for an anti derivative is zero

OpenStudy (anonymous):

ok, now that i have finished my rant, it is clear how to go from \[\ln(\sec(x)+\tan(x)\] to \[\ln(v+\sqrt{1+v^2})\]?

OpenStudy (unklerhaukus):

all this confusion and trig identities AGH i was just trying to get between step in this ODE

OpenStudy (unklerhaukus):

it is so trivial, and i am sorry i asked

OpenStudy (anonymous):

yeah they are just stating it as if it is known that is all, but as you can see the step is either trivial if you look in the back of the book or amazingly annoying if you do it by hand

OpenStudy (unklerhaukus):

i wrote that by the way

OpenStudy (anonymous):

oh yes, integral 25!!

OpenStudy (unklerhaukus):

yeah int # [25]

OpenStudy (unklerhaukus):

i am glad that there are some people out there that can fill in the details when necessary

OpenStudy (unklerhaukus):

thanks